How long after Car B started the race wil Car B catch up with Car A? Express the time in terms of given quantities. Review Cars A and B are racing each other along the same straight road in the following manner: Car A has a head start and is a distance DA beyond the starting line at t 0. The starting line is at z-0. Car A travels at a constant speed vA Car B starts at the starting line but has a better engine than Car A, and thus Car B travels at a constant speed vp, which is greater than vA View Available Hint(s) catch Submit ▼ Part B How far from Car Bs starting line will the cars be when Car B passes Car A? Express your answer in terms of known quantities. (You may use tcatch as well.) View Available Hint(s)
Answer #1

Part A.

Given that Initial distance between Car A and Car B is DA, (where Car A is ahead of car B) And

Car A is moving at a constant speed of vA

Car B is moving at a constant speed of vB

Now at the moment when car B catches car A, then

distance traveled by car = distance between car A and Car B + distance traveled by car A

dB = DA + dA

Now we know that distance = Speed*time, Suppose car B catch car A in 'tcatch' time, then in this time

dB = vB*tcatch

dA = vA*tcatch

So,

vB*tcatch = DA + vA*tcatch

tcatch*(vB - vA) = DA

tcatch = DA/(vB - vA)

DA tcatch = VB - VA

Part B.

when the car B passes car A, then distance from Car B's starting line will be dB, So from above we know that

dB = vB*tcatch

dB = vB*DA/(vB - vA)

dg = VB*DA VB – VA